OpenStudy (keigh2015):

Simplify (3xy2)(4xy)(2xy)3. 24x3y2 24x5y6 96x5y6 96x3y3

OpenStudy (keigh2015):

@undeadknight26 @poppies @TillLindemann

OpenStudy (keigh2015):

@alexistheking777

OpenStudy (anonymous):

three at the end included?

OpenStudy (keigh2015):

it is an exponent

OpenStudy (anonymous):

What is 3xy^2 * 4xy? Don't forget like terms.

OpenStudy (keigh2015):

24xy

OpenStudy (anonymous):

It is 12x^2y^3 because 4*3=12, x * x is X^1+1 which is x^2 and y^2 * y is y^2+1 which is y^3. When yout put it all together it is 12x^2y^3. Now what is that times (2xy)^3/

OpenStudy (keigh2015):

24x^5y^6

OpenStudy (keigh2015):

So the answer is b?

OpenStudy (anonymous):

No what did you forget?

OpenStudy (keigh2015):

The like terms?

OpenStudy (anonymous):

I'm kinda lost with the last part hold on @Calvin3 @AWillie98 @DullJackel09

OpenStudy (anonymous):

c. (3 x y^2) (4 x y) (2 x y)^3\[\left(3 x y^2\right) (4 x y) \left(8 x^3 y^3\right) \]\[\left(12 x^2 y^3\right) \left(8 x^3 y^3\right) \]\[96 x^5 y^6 \]

OpenStudy (keigh2015):

Oh that makes sense

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (keigh2015):

your welcome